# Phase Transitions in Nonreciprocal Driven-Dissipative Condensates

### v2

@@ Abstract
We investigate the influence of boundaries and spatial nonreciprocity on nonequilibrium driven-dissipative phase transitions. We focus on a one-dimensional lattice of nonlinear bosons described by a Lindblad master equation, where the interplay between coherent and incoherent dynamics generates nonreciprocal interactions between sites. Using a mean-field approach, we analyze the phase diagram under both periodic and open boundary conditions. For periodic boundaries, the system always forms a condensate at nonzero momentum and frequency, resulting in a time-dependent traveling wave pattern. In contrast, open boundaries reveal a far richer phase diagram, featuring multiple static and dynamical phases, as well as exotic phase transitions, including the spontaneous breaking of particle-hole symmetry associated with a critical exceptional point and phases with distinct bulk and edge behavior. Our model avoids post-selection or unphysical non-Hermitian Hamiltonians and is experimentally realizable in platforms such as superconducting circuits.
@@
Another example of nonreciprocity arises when a system exhibits preferred spatial directionality [28, 29, 30, 31, 32, 33]. A paradigmatic example is the Hatano-Nelson model [34, 35], a one-dimensional system governed by a non-Hermitian Hamiltonian with asymmetric left and right hopping rates. This asymmetry causes exponential localization of all eigenstates at one boundary under open boundary conditions (OBC), while under periodic boundary conditions (PBC), the eigenstates remain extended—a phenomenon known as the non-Hermitian skin effect (NHSE) [36, 37, 38]. While recent studies have explored nonlinear extensions of such models [39, 40, 41, 42, 43, 44, 31, 27], the impact on phase transitions and the specific role of boundaries, particularly beyond the framework of non-Hermitian Hamiltonians, remains largely underexplored.
@@
In this work, we investigate the effect of nonreciprocity and the NHSE in the presence of interactions on nonequilibrium phase transitions. We focus on a mean-field treatment of driven-dissipative condensates in one dimension, described by a Lindblad master equation, and find that the steady-state phase diagram is highly sensitive to boundary conditions. While periodic boundary conditions lead to a single phase characterized by a finite momentum and frequency condensate, open boundaries result in a rich variety of static and dynamical phases. This includes exotic phase transitions associated with particle-hole symmetry breaking and critical exceptional points (CEPs) [45, 46, 47, 24, 48, 49, 25, 12, 13, 9], as well as phases with distinct bulk and edge dynamics.
@@
The quadratic system ($\Gamma=0$) can be dynamically unstable; with these interaction terms ($\Gamma\neq 0$) the system is always stable and becomes genuinely many-body. Here, we study the mean-field dynamics of the Lindbladian, described by the equation of motion (EOM)
$i\partial_{t}\alpha_{j}$ $=i(\kappa-2\gamma)\alpha_{j}-i\Gamma\absolutevalue{\alpha_{j}}^{2}\alpha_{j}$ (2)
@@
where $\alpha_{j}\equiv\expectationvalue{a_{j}}$ is the coherent state amplitude on each site. In the reciprocal case, $\theta=-\pi/2$, Eq. 2 is equivalent to a discrete Gross-Pitaevskii equation, which is frequently used in the study of driven-dissipative condensates [54, 55, 56, 57, 58].
@@
Periodic boundary conditions.—We begin with a study of the steady-state solutions of Eq. 2 with PBC, i.e., $\alpha_{N+1}=\alpha_{1}$. For large $N$, the vacuum state $\alpha_{j}=0$ is unstable for any $\kappa>0$. This can be seen from the spectrum of Eq. 1 which is diagonalized in momentum space $H_{\mathrm{HN}}=\sum_{q}(i\kappa-i\gamma_{q}+\omega_{q})\outerproduct{q}{q}$ | where
@@
Open boundary conditions.—For PBC, the angle $\theta$ (and hence the nonreciprocity $|J_{+}|-|J_{-}|$) does not play a significant role. In contrast, for OBC, the value of $\theta$ is crucial to the physics. At $\theta=0,\pi$, the hopping is maximally nonreciprocal, and the full nonlinear system has an antilinear $\mathbb{Z}_{2}$ particle-hole $\mathcal{PH}$ symmetry. At the level of the EOM, Eq. 2, this symmetry implies that if $\alpha_{j}(t)$ is a solution, then so is $\mathcal{PH}[\alpha_{j}(t)]=e^{i\pi j}\alpha_{j}^{*}(t)$. We will demonstrate below that this symmetry can be spontaneously broken for OBC. At the level of the linear $\Gamma=0$ theory, this symmetry transforms the quadratic Hatano-Nelson Hamiltonian as $(\mathcal{PH})H_{\mathrm{HN}}(\mathcal{PH})^{-1}=-H_{\mathrm{HN}}$ [62]. In contrast, for more general $\theta\neq 0,\pi$, the $\mathcal{PH}$ symmetry is explicitly broken, making nontrivial steady states time-dependent. Here, we focus on the $\mathcal{PH}$-symmetric, maximally nonreciprocal case of $\theta=0,\pi$, choosing $\theta=\pi$ for concreteness. The corresponding phase diagram is shown in Fig. 3(a).
@@
For $\gamma>J$, as $\kappa$ is increased above $\kappa_{\mathrm{crit}}$, the system undergoes a continuous phase transition from the vacuum phase ($\expectationvalue{\alpha_{j}}_{j}=0$) to a time-independent static condensate [blue region in Fig. 3(a)] with $\expectationvalue{\alpha_{j}}_{j}\neq 0$ [Fig. 3(c)], where $\expectationvalue{o}_{j}$ denotes the spatial average of $o$. Up to a spontaneously chosen global $U(1)$ phase, the complex amplitudes in the static phase are given by $\alpha_{j}^{s}=r_{j}^{s}e^{iq_{s}j}$, where $q_{s}=\pi/2$ ($-\pi/2$) for $\theta=\pi$ ($\theta=0$). Interestingly, this phase order corresponds to the minimal-damping wavevector of the PBC system in Eq. 3. This finite-momentum condensate emerges as $\kappa$ increases in the form of a kink that sweeps across the system from right to left [Fig. 3(b)]. Numerically, we find that the height of the kink is $\sqrt{\kappa/\Gamma}$. In fact, the uniform state $\alpha_{j}=\sqrt{\kappa/\Gamma}e^{i\pi j/2}$ solves Eq. 2 (with $\theta=\pi$) for all $j$ except at the edges ($j=1,N$). As $\kappa\to\infty$, the system approaches this uniform state.
@@
Decreasing $\kappa$ along the leftmost white dashed line in Fig. 3(a) towards the red line, the dynamical matrix describing fluctuations of the static condensate develops a second zero eigenvalue, whose eigenvector coalesces with the Goldstone mode (associated with the spontaneous $U(1)$ symmetry breaking). We thus identify this portion of the red line as a second-order CEP [59]. Below the red line of Fig. 3(a), the system spontaneously breaks the $\mathcal{PH}$ symmetry, giving rise to two bistable time-dependent traveling wave states (wavevectors $q_{\pm}=q_{s}\pm\epsilon$) analogous to the PBC solutions in Eq. 4. The wavevectors $q_{\pm}$ evolves continuously as $\kappa$ is lowered, bifurcating from $q_{\pm}=q_{s}=\pi/2$ in the static phase ($\epsilon=0$) to either $q_{-}=0$ or $q_{+}=\pi$ at the point where the system enters the vacuum phase ($\epsilon=\pi/2$). Similarly, as shown in Fig. 3(d), the frequency evolves continuously from zero, bifurcating into two branches corresponding to clockwise (CW) or counterclockwise (CCW) rotations (analogous to the “chiral” states in Ref. [8]). The $\mathcal{PH}$ symmetry maps one solution onto the other, transforming a CW rotating state into a CCW one (i.e., flipping the sign of $\omega$), while simultaneously flipping the wavevector from $q_{s}+\epsilon$ to $q_{s}-\epsilon$, and vice versa. Remarkably, the PBC result for the frequency in Eq. 3 holds perfectly for these two states, using the numerically extracted average wavevector $\langle q\rangle_{j}$.
@@
Edge and bulk phase transitions.— So far, we have considered cases where only the phase of each site varied in time. However, in certain regions of the dynamical phase, the amplitudes also fluctuate. These fluctuations occur predominantly at the left edge, leading to distinct bulk versus edge behavior [Fig. 4(a)]. In fact, Fig. 3(d) already contains such an example. This is illustrated in Fig. 3(e), which displays the time- and space-averaged absolute value of the rate of change of the magnitude, $\langle|\partial_{t}r_{j}^{2}|\rangle_{j,t}$. This quantity becomes nonzero at $\kappa/J\sim 1.9-2.2$. The inset in Fig. 3(e) shows an example of the time-averaged spatial profile ($\expectationvalue{|\partial_{t}r_{j}^{2}|}_{t}$) of these amplitude fluctuations, revealing that they are peaked at the left edge. Away from the edge, this quantity quickly drops, and then exponentially decays into the bulk. Notably, these amplitude fluctuations do not affect the $\langle\omega\rangle_{j}$ order parameter shown in Fig. 3(d), as the dominant Fourier component, even for the left-most sites, remains at the same bulk value as the average value $\langle\omega\rangle_{j}$ depicted in Fig. 3(d).
The dynamical region ($\partial_{t}\alpha_{j}\neq 0$) includes additional phases with distinct edge and bulk behaviors, with some regions exhibiting multistability, indicated by stripes in Fig. 4(a). Phase IV, indicated in purple in Fig. 4(a), features amplitude fluctuations at the left edge and phase-only oscillations in the bulk. A representative example ($\gamma=0.3,\kappa=2.2$) of the amplitude and phase space-time dynamics is shown in Fig. 4(b,i) and Fig. 4(b,ii), respectively. Left of the black dashed line in Fig. 4(a), the two limit cycles of Phase I – the CW and the CCW solutions – merge, resulting in more complex dynamics. The phases of the bulk sites act as pendula, involving both CW and CCW rotations [see Fig. 4(e)]. In some parts of this phase, particularly at smaller $\gamma$, the dynamics is nevertheless periodic. Notably, the $\mathbb{Z}_{2}$ symmetry is dynamically restored in this case, corresponding to $\mathcal{PH}$ accompanied by the breaking of continuous time translation symmetry [63, 64]. Specifically, the $\mathcal{PH}$ operation acts as a translation by a half period, and the steady state satisfies $\alpha_{j}(t)=\mathcal{PH}[\alpha_{j}(t+T/2)]$ for some period $T$ [59].

### v3

@@ Abstract
We investigate the influence of boundaries and spatial nonreciprocity on nonequilibrium driven-dissipative phase transitions. We focus on a one-dimensional lattice of nonlinear bosons described by a Lindblad master equation, where the interplay between coherent and incoherent dynamics generates nonreciprocal interactions between sites. Using a mean-field approach, we analyze the phase diagram under both periodic and open boundary conditions. For periodic boundaries, the system always forms a condensate at nonzero momentum and frequency, resulting in a time-dependent traveling wave pattern. In contrast, open boundaries reveal a far richer phase diagram, featuring multiple static and dynamical phases, as well as exotic phase transitions, including the spontaneous breaking of particle-hole symmetry associated with a critical exceptional point and phases with distinct bulk and edge behavior. Our model does not require post-selection and is experimentally realizable in platforms such as superconducting circuits.
@@
Another example of nonreciprocity arises when a system exhibits preferred spatial directionality [28, 29, 30, 31, 32, 33]. A paradigmatic example is the Hatano-Nelson model [34, 35], a one-dimensional system governed by a non-Hermitian Hamiltonian with asymmetric left and right hopping rates. This asymmetry causes exponential localization of all eigenstates at one boundary under open boundary conditions (OBC), while under periodic boundary conditions (PBC), the eigenstates remain extended—a phenomenon known as the non-Hermitian skin effect (NHSE) [36, 37, 38]. While recent studies have explored nonlinear extensions of such models [39, 40, 41, 42, 43, 44, 31, 27, 45], the impact on phase transitions and the specific role of boundaries, particularly beyond the framework of non-Hermitian Hamiltonians, remains largely underexplored.
@@
In this work, we investigate the interplay of nonreciprocity, interactions, and boundary conditions on nonequilibrium phase transitions in a minimal driven-dissipative condensate. Driven dissipative condensates have been extensively studied across a range of systems [46, 47, 48, 49, 50, 2, 51, 52, 53, 54, 55, 56]. Recent theoretical [57] and experimental [58] works have begun to explore condensates in nonreciprocal settings, but the impact of nonreciprocity on phase transitions remains unexplored. We focus on a mean-field treatment of a one dimensional condensate described by a Lindblad master equation, and find that the steady-state is highly sensitive to boundary conditions. While periodic boundary conditions lead to a single phase characterized by a finite momentum and frequency condensate, open boundaries result in a rich variety of static and dynamical phases. These include exotic phase transitions associated with the breaking and dynamical restoration of particle-hole symmetry, boundary-induced critical exceptional points [59, 60, 61, 24, 62, 63, 25, 12, 13, 9], and regimes with qualitatively distinct edge and bulk dynamics.
@@
The quadratic system ($\Gamma=0$) can be dynamically unstable; with these interaction terms ($\Gamma\neq 0$) the system is always stable and becomes genuinely many-body. Here, we study the mean-field dynamics of the Lindbladian, described by the equation of motion
[…]
@@
where $\alpha_{j}\equiv\expectationvalue{\hat{a}_{j}}$ is the coherent state amplitude on each site.
@@
Periodic boundary conditions.—We begin with a study of the steady-state solutions of Eq. 2 with PBC, i.e., $\alpha_{N+1}=\alpha_{1}$. For large $N$, the vacuum state $\alpha_{j}=0$ is unstable for any $\kappa>0$. This can be seen from the spectrum of Eq. 1 which is diagonalized in momentum space $|H_{\mathrm{HN}}=\sum_{q}(i\kappa-i\gamma_{q}+\omega_{q})\outerproduct{q}{q}$ where
@@
Open boundary conditions.—For PBC, the angle $\theta$ (and hence the nonreciprocity $|J_{+}|-|J_{-}|$) does not play a significant role. In contrast, for OBC, the value of $\theta$ is crucial to the physics. At $\theta=0,\pi$, the hopping is maximally nonreciprocal, and the full nonlinear system has an antilinear $\mathbb{Z}_{2}$ particle-hole $\mathcal{PH}$ symmetry. At the level of Eq. 2, this symmetry implies that if $\alpha_{j}(t)$ is a solution, then so is $\mathcal{PH}[\alpha_{j}(t)]=e^{i\pi j}\alpha_{j}^{*}(t)$. We will demonstrate below that this symmetry can be spontaneously broken for OBC. At the level of the linear $\Gamma=0$ theory, this symmetry transforms the quadratic Hatano-Nelson Hamiltonian as $(\mathcal{PH})H_{\mathrm{HN}}(\mathcal{PH})^{-1}=-H_{\mathrm{HN}}$ [71]. In contrast, for more general $\theta\neq 0,\pi$, the $\mathcal{PH}$ symmetry is explicitly broken, making nontrivial steady states time-dependent. Here, we focus on the $\mathcal{PH}$-symmetric, maximally nonreciprocal case of $\theta=0,\pi$, choosing $\theta=\pi$ for concreteness. The corresponding phase diagram is shown in Fig. 3(a).
@@
For $\gamma>J$, as $\kappa$ is increased above $\kappa_{\mathrm{crit}}$, the system undergoes a continuous phase transition from the vacuum phase ($\expectationvalue{\alpha_{j}}_{j}=0$) to a time-independent static condensate [blue region in Fig. 3(a)] with $\expectationvalue{\alpha_{j}}_{j}\neq 0$ [Fig. 3(c)], where $\expectationvalue{o}_{j}$ denotes the spatial average of $o$. Up to a spontaneously chosen global $U(1)$ phase, the complex amplitudes in the static phase are given by $\alpha_{j}^{s}=r_{j}^{s}e^{iq_{s}j}$, where $q_{s}=\pi/2$ ($\pi/2$) for $\theta=\pi$ ($\theta=0$). Interestingly, this phase order corresponds to the minimal-damping wavevector of the PBC system in Eq. 3. This finite-momentum condensate emerges as $\kappa$ increases in the form of a kink that sweeps across the system from right to left [Fig. 3(b)]. Numerically, we find that the height of the kink is $\sqrt{\kappa/\Gamma}$. In fact, the uniform state $\alpha_{j}=\sqrt{\kappa/\Gamma}e^{i\pi j/2}$ solves Eq. 2 (with $\theta=\pi$) for all $j$ except at the edges ($j=1,N$). As $\kappa\to\infty$, the system approaches this uniform state.
@@
Decreasing $\kappa$ along the leftmost white dashed line in Fig. 3(a) towards the red line, the dynamical matrix describing fluctuations of the static condensate develops a second zero eigenvalue, whose eigenvector coalesces with the Goldstone mode (associated with the spontaneous $U(1)$ symmetry breaking). We thus identify this portion of the red line as a second-order critical exceptional point [68]. Below the red line of Fig. 3(a), the system spontaneously breaks the $\mathcal{PH}$ symmetry, giving rise to two bistable time-dependent traveling wave states (wavevectors $q_{\pm}=q_{s}\pm\epsilon$) analogous to the PBC solutions in Eq. 4. The wavevectors $q_{\pm}$ evolves continuously as $\kappa$ is lowered, bifurcating from $q_{\pm}=q_{s}=\pi/2$ in the static phase ($\epsilon=0$) to either $q_{-}=0$ or $q_{+}=\pi$ at the point where the system enters the vacuum phase ($\epsilon=\pi/2$). Similarly, as shown in Fig. 3(d), the frequency evolves continuously from zero, bifurcating into two branches corresponding to clockwise (CW) or counterclockwise (CCW) rotations (analogous to the “chiral” states in Ref. [8]). The $\mathcal{PH}$ symmetry maps one solution onto the other, transforming a CW rotating state into a CCW one (i.e., flipping the sign of $\omega$), while simultaneously flipping the wavevector from $q_{s}+\epsilon$ to $q_{s}-\epsilon$, and vice versa. Remarkably, the PBC result for the frequency in Eq. 3 holds perfectly for these two states, using the numerically extracted average wavevector $\langle q\rangle_{j}$.
@@
Edge and bulk phase transitions.— So far, we have considered cases where only the phase of each site varied in time. In this section, we will briefly discuss some additional dynamical phases where the amplitudes also fluctuate. These fluctuations occur predominantly at the left edge, leading to distinct bulk versus edge behavior [Fig. 4(a)]. In fact, Fig. 3(d) already contains such an example. This is illustrated in Fig. 3(e), which displays the time- and space-averaged absolute value of the rate of change of the magnitude, $\langle|\partial_{t}r_{j}^{2}|\rangle_{j,t}$. This quantity becomes nonzero at $\kappa/J\sim 1.9-2.2$. The inset in Fig. 3(e) shows an example of the time-averaged spatial profile ($\expectationvalue{|\partial_{t}r_{j}^{2}|}_{t}$) of these amplitude fluctuations, revealing that they are peaked at the left edge. Away from the edge, this quantity quickly drops, and then exponentially decays into the bulk. Notably, these amplitude fluctuations do not affect the $\langle\omega\rangle_{j}$ order parameter shown in Fig. 3(d), as the dominant Fourier component, even for the left-most sites, remains at the same bulk value as the average value $\langle\omega\rangle_{j}$ depicted in Fig. 3(d).
The dynamical region ($\partial_{t}\alpha_{j}\neq 0$) includes additional phases with distinct edge and bulk behaviors, with some regions exhibiting multistability, indicated by stripes in Fig. 4(a). Phase IV, indicated in purple in Fig. 4(a), features amplitude fluctuations at the left edge and phase-only oscillations in the bulk. A representative example ($\gamma=0.3,\kappa=2.2$) of the amplitude and phase space-time dynamics is shown in Fig. 4(b,i) and Fig. 4(b,ii), respectively. Left of the black dashed line in Fig. 4(a), the two limit cycles of Phase I – the CW and the CCW solutions – merge, resulting in more complex dynamics. The phases of the bulk sites act as pendula, involving both CW and CCW rotations [see Fig. 4(e)]. In some parts of this phase, particularly at smaller $\gamma$, the dynamics is nevertheless periodic. Notably, the $\mathbb{Z}_{2}$ symmetry is dynamically restored in this case, corresponding to $\mathcal{PH}$ accompanied by the breaking of continuous time translation symmetry [72, 73]. Specifically, the $\mathcal{PH}$ operation acts as a translation by a half period, and the steady state satisfies $\alpha_{j}(t)=\mathcal{PH}[\alpha_{j}(t+T/2)]$ for some period $T$ [68]. Additional details on these phases and their symmetry classification are provided in the End Matter.